OpenStudy (anonymous):

? A circle is described by the equation . What are the cordinates for the center of the circle and the length of the radius?

OpenStudy (anonymous):

(-12,-32), 23units (-12,-32), 49units (12,32), 49units (12,32), 23units

OpenStudy (mathstudent55):

Where is the equation?

OpenStudy (anonymous):

(x−12)2+(y−32)2=49.

OpenStudy (anonymous):

thats the equation @mathstudent55

OpenStudy (mathstudent55):

The equation of a circle with center at (h, k) nad radius r is \(\large (x - h)^2 + (y - k)^2 = r^2\) Compare the equation you were given with the form above. What are your values of h, k, and r?

OpenStudy (anonymous):

12, 32, and 49? @mathstudent55

OpenStudy (mathstudent55):

12 and 32 are correct for h and k, respectively, so the center is (12, 32). Be careful with the radius. The right side has to be written as r^2. Then r is the radius. Your right side is 49. How do you write 49 as the square of something?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

h= .25 k=2.25 r= .2?

OpenStudy (mathstudent55):

Correct. \((x−12)^2+(y−32)^2=7^2\) Center: (12, 32) Radius: 7

OpenStudy (mathstudent55):

I have a question for you. In your choices, you have radius only of 49 and 23. Did your equation have 49 or 49^2 on the right side?

OpenStudy (anonymous):

49

OpenStudy (anonymous):

thankyou so much though, i got the right answer :D @mathstudent55

OpenStudy (mathstudent55):

Great.

OpenStudy (mathstudent55):

You're welcome.